OpenStudy (anonymous):

will medal and fan is x - 2 is a factor of p(x) = x^4 - 3x^2 + 2x - 8?

OpenStudy (prince_vegeta):

yes

OpenStudy (prince_vegeta):

p(x)=(x−2)(x+2.314596)

OpenStudy (prince_vegeta):

so yes

OpenStudy (jadedry):

There is a neat trick you can use in order to find out whether something is a factor or not. for a function f(x) , and a factor (x+a) The remainder is f(-a) , if the remainder ends up being zero, then (x+a) is a factor. as we have (x-2) , we have to plug -(-2) into the function. So we have: \[p(2) = 2^4 - 3(2^2) + (2*2) - 8\] =\[p(2) = 16 - 12 + 4 - 8 = 0\] x-2 is a factor! c:

OpenStudy (prince_vegeta):

thanks for proving my point

OpenStudy (anonymous):

thank you!!